If f(x) = log 5x, what is f(1/5)
Answer by drk(1908)

If f(x) = log_5x, what is f(1/5)?
THe x goes away and the number comes to stay.
F(1/5) = log_5(1/5)
What is the value of log_5(1/5)?
5^y = 1/5
y = -1